# Optimal Client Sampling in Federated Learning with Client-Level Heterogeneous Differential Privacy

### v1

@@ B.4 Additional Empirical Results
Impact of Clipping Threshold on Model Utility. We now analyze the influence of the clipping threshold $C$ on model utility. Increasing $C$ results in less aggressive clipping of local model updates, but also amplifies the magnitude of the noise required to satisfy differential privacy constraints.
We evaluate the impact of larger clipping thresholds by scaling $C$ using multiplicative factors: 1.25×1.25$\times$1.25 × and 1.5×1.5$\times$1.5 ×. The results, presented in Table 10, indicate that as $C$ increases, the performance of DP-FedAvg degrades significantly, particularly on FMNIST and SVHN datasets. In contrast, our method, GDPFed+, exhibits only minor fluctuations in performance under each setting, demonstrating strong robustness. Notably, on SVHN, GDPFed+ achieves an average test accuracy of $70.26\%$, representing a substantial improvement of $+37.39\%$ over DP-FedAvg. This performance gain is largely attributed to the use of optimized client sampling ratios, which yield more favorable noise multipliers for each privacy group. These results further underscore the effectiveness and practical resilience of GDPFed+ under varying clipping thresholds.
[…]
Dataset Method 1.0×1.0$\times$1.0 × 1.25×1.25$\times$1.25 × 1.5×1.5$\times$1.5 × Average
[…]
B.5 Broader Impacts
This work addresses the challenge of client-level HDPFL, where individual clients have their privacy preferences. While our proposed method, GDPFed+, effectively reduces the noise added to the global model and improves utility under strict privacy constraints, it assumes that clients’ privacy budgets are externally specified and fixed. In real-world deployments, however, determining an adequate privacy level $\epsilon_{i}$ for each individual $i\in[n]$ is non-trivial and often subject to misunderstanding or misuse. To ensure the ethical application of our method, the assignment of individual privacy guarantees must be transparent, informed, and not subject to coercion.
While our work improves the privacy-utility trade-off in HDPFL, it also highlights the importance of coupling technical advances with social mechanisms to promote responsible, equitable, and informed use of DP in practice.
Appendix C Proof of Theorem 1
Credits: our proof follows the proof of Theorem 1 in [25].
Proof.
Suppose the client is sampled without replacement with probability $q_{m}$ at each round for each group $m\in[M]$. By Lemma 6 and Lemma 7, the $t$-th round of GDPFed satisfies $(\alpha_{m},\rho_{t}(\alpha_{m}))$-RDP for each group $m$, where
$\rho_{t}(\alpha)=\frac{3.5q_{m}^{2}\alpha_{m}}{\sigma_{m}^{2}},$ (1)
if $\sigma_{m}^{2}\geq 0.7$ and $\alpha_{m}\leq 1+(2/3)C^{2}\sigma^{2}_{m}\log(1/q_{m}\alpha_{m}(1+\sigma^{2}_{m}))$. Then by Lemma 5, each group in GDPFed satisfies $(\alpha_{m},T\rho_{t}(\alpha_{m}))$-RDP after $T$ rounds of training. Next, in order to guarantee $(\epsilon_{m},\delta)$-DP according to Lemma 4, we need
$\frac{3.5q^{2}_{m}T\alpha_{m}}{\sigma^{2}_{m}}+\frac{\log(1/\delta)}{\alpha_{m}-1}\leq\epsilon_{m}.$ (2)
Suppose $\alpha_{m}$ and $\sigma_{m}$ are chosen such that the conditions for Equation (1) are satisfied. Choose $\alpha_{m}=1+2\log(1/\delta_{m})/\epsilon_{m}$ and rearrange the inequality in Inequality (2), we need
$\sigma^{2}_{m}\geq\frac{7q^{2}_{m}T(\epsilon_{m}+2\log(1/\delta))}{\epsilon^{2}_{m}}.$ (3)
Then, using the constraint on $\epsilon$ concludes the proof. ∎
Appendix D Proof of Convergence Bound of Sparsification-Amplified GDPFed
D.1 Useful Lemmas
Lemma 8.
Given any two vectors $a,b\in\mathbb{R}^{d}$,
$0\displaystyle 2\left<a,b\right>\leq\alpha\left\|a\right\|^{2}+\frac{1}{\alpha}\left\|b\right\|^{2},\forall\;\alpha>$ 0 .
Lemma 9.
Given any two vectors $a,b\in\mathbb{R}^{d}$,
$\left\|a+b\right\|^{2}\leq(1+\delta)\left\|a\right\|^{2}+(1+\delta^{-1})\left\|b\right\|^{2},\forall\;\delta>0.$
Lemma 10.
Given arbitrary set of n vectors $\{a_{i}\}^{n}_{i=1}$, $a_{i}\in\mathbb{R}^{d}$,
$\left\|\sum^{n}_{i=1}a_{i}\right\|^{2}\leq n\sum^{n}_{i=1}\left\|a_{i}\right\|^{2}.$
D.2 Proof of Theorem 3
Notations. We let $\nabla f_{m,i}({\theta}_{m,i}^{t,s})$ represent the local gradient for client $i$ in group $m$ so that $\mathbb{E}[\mathbf{g}_{m,i}^{t,s}]=\nabla f_{m,i}(\theta_{m,i}^{t,s})$. For ease of expression, we let $\mathbf{d}^{t}_{m,i}=(1/\tau)\sum_{s=0}^{\tau-1}\mathbf{g}_{m,i}^{t,s}$ and $\mathbf{h}_{m,i}^{t}=(1/\tau)\sum_{s=0}^{\tau-1}\nabla f_{m,i}(\theta_{m,i}^{t,s})$, and $\mathbb{E}_{t}[\mathbf{d}_{i}^{t}]=\mathbf{h}_{i}^{t}$. We have the update rule $\theta^{t+1}=\theta^{t}-(1/\sum^{M}_{j=1}r_{j})\sum_{m=1}^{M}\omega_{m}(\mathbf{y}^{t}_{m}\odot\mathbf{mk}^{t}_{m})$, where $\omega_{m}$ is a reweighting parameter and $\mathbf{y}^{t}_{m}=\sum_{i\in\mathcal{S}^{t}_{m}}\Delta^{t}_{m,i}$. More specifically:
$\frac{1}{\sum^{M}_{j=1}r_{j}}\sum_{m=1}^{M}\omega_{m}(\mathbf{y}^{t}_{m}\odot\mathbf{mk}^{t}_{m})$ $=\sum_{m=1}^{M}\frac{1}{\sum^{M}_{j=1}r_{j}}\omega_{m}(\mathbf{y}^{t}_{m}\odot\mathbf{mk}^{t}_{m})$
$\leq\sum_{m=1}^{M}\frac{1}{r_{m}}\omega_{m}(\mathbf{y}^{t}_{m}\odot\mathbf{mk}^{t}_{m})$
$=\sum^{M}_{m=1}\omega_{m}\frac{1}{|\mathcal{S}_{m}|}\left(\sum_{i\in\mathcal{S}_{m}}\left(\eta\tau\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})+\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,i})\right)\right)$
$=\sum^{M}_{m=1}\omega_{m}\frac{1}{|\mathcal{S}_{m}|}\left(\sum_{i\in\mathcal{S}_{m}}\eta\tau\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})+\sum_{i\in\mathcal{S}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,i})\right)$
$=\sum^{M}_{m=1}\omega_{m}\frac{1}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\eta\tau\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})+\sum^{M}_{m=1}\omega_{m}\frac{1}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,i}).$
Proof.
With Assumption 1, we have
$f(\theta^{t+1})-f(\theta^{t})$ $\leq\mathbb{E}_{t}\langle\nabla f(\theta^{t}),\theta^{t+1}-\theta^{t}\rangle+\frac{L}{2}\mathbb{E}_{t}\|\theta^{t+1}-\theta^{t}\|^{2}$
$\leq\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),-\sum_{m=1}^{M}\omega_{m}\tilde{\Delta}_{m}^{t}\right\rangle+\frac{L}{2}\mathbb{E}_{t}\|\theta^{t+1}-\theta^{t}\|^{2}$
$=\underbrace{-\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\omega_{m}\left[\frac{1}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\eta\tau\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})+\frac{1}{|\mathcal{S}_{m}|}\sum_{j\in\mathcal{S}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,j})\right]\right\rangle}_{T_{1}}$
$\quad+\underbrace{\frac{L}{2}\mathbb{E}_{t}\|\theta^{t+1}-\theta^{t}\|^{2}}_{T_{2}}.$
To solve $T_{1}$, we have
$T_{1}$ $=-\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\eta\tau\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})\right)\right\rangle-\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,i})\right\rangle$
$=-\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\eta\tau\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})+\eta\tau\mathbf{d}^{t}_{m,i}-\eta\tau\mathbf{d}^{t}_{m,i}\right)\right\rangle$
$=-\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\eta\tau\mathbf{d}^{t}_{m,i}\right)\right\rangle+\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\eta\tau\mathbf{d}^{t}_{m,i}-\eta\tau\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})\right)\right\rangle$
$=-\eta\tau\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\omega_{m}\mathbb{E}_{t}\left[\frac{1}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\mathbf{d}^{t}_{m,i}\right]\right\rangle+\eta\tau\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\mathbf{d}^{t}_{m,i}-\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})\right)\right\rangle$
$=\underbrace{-\eta\tau\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\mathbf{h}^{t}_{m,i}\right\rangle}_{A_{1}}+\underbrace{\eta\tau\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\mathbf{d}^{t}_{m,i}-\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})\right)\right\rangle}_{A_{2}}.$
For $A_{1}$, we have
$A_{1}$ $=-\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})\right\|^{2}-\frac{\eta\tau}{2}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\mathbf{h}^{t}_{m,i}\right\|^{2}+\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})-\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\mathbf{h}^{t}_{m,i}\right\|^{2}$
$=-\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})-\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\frac{1}{\tau}\sum^{\tau-1}_{s=0}\nabla f_{m,i}(\theta^{t,s}_{m,i})\right\|^{2}$
$=-\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau}{2}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\nabla f_{m,i}(\theta^{t})-\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\frac{1}{\tau}\sum^{\tau-1}_{s=0}\nabla f_{m,i}(\theta^{t,s}_{m,i})\right\|^{2}$
$=-\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau}{2}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\frac{1}{\tau}\sum^{\tau-1}_{s=0}\left(\nabla f_{m,i}(\theta^{t})-\nabla f_{m,i}(\theta^{t,s}_{m,i})\right)\right\|^{2}$
$\leq-\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau}{2}\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\frac{1}{\tau}\sum^{\tau-1}_{s=0}\left\|\nabla f_{m,i}(\theta^{t})-\nabla f_{m,i}(\theta^{t,s}_{m,i})\right\|^{2}$
$\leq-\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau L^{2}}{2}\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\frac{1}{\tau}\sum^{\tau-1}_{s=0}\left\|\theta^{t}-\theta^{t,s}_{m,i}\right\|^{2},$
where the first inequality holds according to Lemma 10 and the last inequality holds according to Assumption 1.
For $A_{2}$, we have
$A_{2}$ $=\eta\tau\mathbb{E}_{t}\left\langle\nabla f(\theta^{t}),\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\mathbf{d}^{t}_{m,i}-\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})\right)\right\rangle$
$\leq\frac{\eta\tau}{2}\mathbb{E}_{t}\left(\gamma\left\|\nabla f(\theta^{t})\right\|^{2}+\gamma^{-1}\mathbb{E}_{t}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\mathbf{d}^{t}_{m,i}-\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})\right)\right\|^{2}\right)$
$\leq\frac{\eta\tau}{2}\left(\gamma\left\|\nabla f(\theta^{t})\right\|^{2}+\gamma^{-1}\mathbb{E}_{t}\left[\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left\|\mathbf{d}^{t}_{m,i}-\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})\right\|^{2}\right]\right)$
$\leq\frac{\eta\tau\gamma}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau}{2\gamma}\mathbb{E}_{t}\left[\sum_{m=1}^{M}\frac{\omega_{m}\phi_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left\|\mathbf{d}^{t}_{m,i}\right\|^{2}\right]$
$\leq\frac{\eta\tau\gamma}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau}{2\gamma}\sum_{m=1}^{M}\frac{\omega_{m}\phi_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\mathbb{E}_{t}\left\|\mathbf{d}^{t}_{m,i}\right\|^{2},$
where the first inequality follows Lemma 8, the second inequality follows Lemma 10, and the third inequality holds with bounded sparsification (Lemma 3) and $\phi_{m}=\left(1-k/d\right)^{2}$.
Plugging $A_{1},A_{2}$ back to $T_{1}$, we have
$T_{1}$ $\leq-\frac{\eta\tau}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau L^{2}}{2}\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\frac{1}{\tau}\sum^{\tau-1}_{s=0}\left\|\theta^{t}-\theta^{t,s}_{m,i}\right\|^{2}+\frac{\eta\tau\gamma}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau}{2\gamma}\sum_{m=1}^{M}\frac{\omega_{m}\phi_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\mathbb{E}_{t}\left\|\mathbf{d}^{t}_{m,i}\right\|^{2}$
$=\frac{\eta\tau(\gamma-1)}{2}\left\|\nabla f(\theta^{t})\right\|^{2}+\frac{\eta\tau L^{2}}{2}\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\frac{1}{\tau}\sum^{\tau-1}_{s=0}\left\|\theta^{t}-\theta^{t,s}_{m,i}\right\|^{2}+\frac{\eta\tau}{2\gamma}\sum_{m=1}^{M}\frac{\omega_{m}\phi_{m}}{|\mathcal{G}_{m}|}\sum_{i\in\mathcal{G}_{m}}\mathbb{E}_{t}\left\|\mathbf{d}^{t}_{m,i}\right\|^{2}.$
For $T_{2}$, we have
$T_{2}$ $=\frac{L}{2}\mathbb{E}_{t}\left\|\sum_{m=1}^{M}\omega_{m}\left[\frac{1}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\eta\tau\mathrm{Top}_{k_{m}}(\mathbf{d}^{t}_{m,i})+\frac{1}{|\mathcal{S}_{m}|}\sum_{j\in\mathcal{G}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,j})\right]\right\|^{2}$
$=\frac{L\eta^{2}\tau^{2}}{2}\mathbb{E}_{t}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left(\mathrm{Top}_{m,k}^{+}\left(\mathbf{d}^{t}_{m,i}\right)-\mathbf{d}^{t}_{m,i}+\mathbf{d}^{t}_{m,i}\right)\right\|^{2}+\frac{L}{2}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{j\in\mathcal{G}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,j})\right\|^{2}$
$\leq\frac{L\eta^{2}\tau^{2}}{2}\mathbb{E}_{t}\left[\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left\|\left(\mathrm{Top}_{k_{m}}\left(\mathbf{d}^{t}_{m,i}\right)-\mathbf{d}^{t}_{m,i}+\mathbf{d}^{t}_{m,i}\right)\right\|^{2}\right]+\frac{L}{2}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{j\in\mathcal{G}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,j})\right\|^{2}$
$\leq L\eta^{2}\tau^{2}\mathbb{E}_{t}\left[\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left\|\left(\mathrm{Top}_{k_{m}}\left(\mathbf{d}^{t}_{m,i}\right)-\mathbf{d}^{t}_{m,i}\right)\right\|^{2}\right]+L\eta^{2}\tau^{2}\mathbb{E}_{t}\left[\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left\|\mathbf{d}^{t}_{m,i}\right\|^{2}\right]$
$\quad\quad+\frac{L}{2}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{j\in\mathcal{G}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,j})\right\|^{2}$
$\leq L\eta^{2}\tau^{2}\mathbb{E}_{t}\left[\sum_{m=1}^{M}\frac{\omega_{m}\phi_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left\|\mathbf{d}^{t}_{m,i}\right\|^{2}\right]+L\eta^{2}\tau^{2}\mathbb{E}_{t}\left[\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left\|\mathbf{d}^{t}_{m,i}\right\|^{2}\right]+\frac{L}{2}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{j\in\mathcal{G}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,j})\right\|^{2}$
$=L\eta^{2}\tau^{2}\mathbb{E}_{t}\left[\sum_{m=1}^{M}\frac{\omega_{m}\left(\phi_{m}+1\right)}{|\mathcal{S}_{m}|}\sum_{i\in\mathcal{S}_{m}}\left\|\mathbf{d}^{t}_{m,i}\right\|^{2}\right]+\frac{L}{2}\left\|\sum_{m=1}^{M}\frac{\omega_{m}}{|\mathcal{S}_{m}|}\sum_{j\in\mathcal{G}_{m}}\mathrm{Top}_{k_{m}}(\mathbf{b}^{t}_{m,j})\right\|^{2},$
where the first inequality follows Lemma 10, the second inequality follows Lemma 9, the third inequality follows Lemma 3.
Combine $T_{1}$ and $T_{2}$, we have
[…]

### v3

@@ VI-C4 Impact of privacy budget
TABLE VIII: Test accuracy ($\%$) of GDPFed+ across datasets under varying sparsification levels; GDPFed-$\mathrm{opc}$ with configuration $(1.0,1.0,1.0)$ serves as the baseline.
Sparsification Level FMNIST SVHN CIFAR-10 Shakespeare Average
$(0.9,0.9,0.9)$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}75.89}_{\pm 0.53}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.24})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}71.04}_{\pm 0.60}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.19})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}38.69}_{\pm 0.33}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.65})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}41.94}_{\pm 0.42}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.26})$ $56.89\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.33})$
$(0.7,0.7,0.7)$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}75.79}_{\pm 0.44}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.14})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}70.79}_{\pm 0.41}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-0.06})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}38.40}_{\pm 0.42}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.36})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}41.90}_{\pm 0.34}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.22})$ $56.72\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.16})$
$(0.5,0.5,0.5)$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}75.37}_{\pm 0.41}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-0.28})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}70.39}_{\pm 0.65}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-0.46})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}38.53}_{\pm 0.74}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+0.49})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}41.40}_{\pm 0.26}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-0.28})$ $56.42\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-0.14})$
$(0.3,0.3,0.3)$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}74.92}_{\pm 0.02}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-0.73})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}68.94}_{\pm 1.29}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-1.91})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}37.50}_{\pm 0.05}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-0.54})$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}40.33}_{\pm 0.73}\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-1.35})$ $55.42\,({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-1.14})$
[…]